Owl-Eye Radiologists (OR) does various types of diagnostic imaging. Radiologists perform tests using sophisticated equipment. OR's management wants to compute the costs of performing tests for two different types of patients: those who are hospitalized (including those in emergency rooms) and those who are not hospitalized but are referred by physicians. The data for June for the two categories of patients follow.

Hospital Patients Other Patients Total
Units (i.e., procedures) produced 680 870 1,550
Equipment-hours used 240 110 350
Direct labor-hours 470 160 630
Direct labor costs $ 37,600 $ 9,600 $ 47,200
Overhead costs $ 47,950

The accountant first assigns overhead costs to two pools: overhead related to equipment-hours and overhead related to labor-hours. The analysis of overhead accounts by the cost accountant follows.

Account Amount Related to:
Utilities $ 4,820 Equipment-hours
Supplies 12,580 Labor-hours
Indirect labor and supervision 18,920 Labor-hours
Equipment depreciation and maintenance 8,140 Equipment-hours
Miscellaneous 3,490 Equipment-hours


Required:

a. Compute the predetermined overhead rates assuming that Owl-Eye Radiologists uses equipment-hours to allocate equipment-related overhead costs and labor-hours to allocate labor-related overhead costs.

b. Compute the total costs of production and the cost per unit for each of the two types of patients undergoing tests in June.
Answer #1

Predetermined Rate = Overhead cost/Total activity of Driver

Equipment Related Costs = (4820+8140+3490)/350

= $47 per equipment hour

Labor rated = (12580+18920)/630

= $50 per labor hour

b.Total cost = Labor cost + overhead cost

Hospital Patients

Total cost = 37,600+47*240+50*470 = $72,380

Cost per unit = 72,380/680 = $106.44

Other patients

Total cost = 9600 + 47*110 + 50*160 = $22,770

Per unit = 22,770/870 = $26.17